OpenStudy (anonymous):

How do I find the zeros and asymptotes for these two trig. functions? :) 1) y=2cot(3x) 2) y=csc(-1/2 (x))

OpenStudy (tkhunny):

You may wish to find zeros and asymptotes of the more familiar reciprocal functions.

OpenStudy (anonymous):

I don't understand what you mean, sorry. please explain

OpenStudy (tkhunny):

\(y = \cos(x)\) This has zeros at \(\pi/2 + k\pi\) for integer values of \(k\). \(y = \cos(x)\) Has no asymptote. Keeping all that in mind, what can we say of \(y = \sec(x)\)? \(y = \sec(x)\) This has asymptotes at \(\pi/2 + k\pi\) for integer values of \(k\). \(y = \sec(x)\) Has no zero.

OpenStudy (anonymous):

thank you. one more question. What is K?

OpenStudy (tkhunny):

It is explained, above. Any integer value.